# LogHD: Robust Compression of Hyperdimensional Classifiers via Logarithmic Class-Axis Reduction

### v1

@@ I Introduction
Machine-learning architectures operating under tight memory, energy, and reliability budgets increasingly depend on cross-layer co-design and in-/near-memory substrates. Recent surveys and system studies underline both the promise and the sensitivity of memory-centric ML to device non-idealities and limited precision, motivating methods that reduce footprint while preserving robustness [lammie2025deep, krestinskaya2024neural, tsai2024architecture, burr2023design, frank2023impact]. Within this context, hyperdimensional computing (HDC)—a family of vector–symbolic architectures (VSA)—is a strong fit: its operations are inherently parallel and memory-centric, and recent cross-layer workflows and macro-level prototypes indicate practical routes to efficient HDC/VSA hardware in embedded and in-/near-memory settings [nayan2025hydra, du2025cross].
HDC encodes data as $D$-dimensional hypervectors and learns via light-weight algebraic operations that are parallel, memory-centric, and empirically tolerant to device noise [kanerva2009hyperdimensional, yun2024neurohash, imani2019framework, zou2021scalable, wang2023disthd, yun2025hyperdimensional, yun2023hyperdimensional, yun2024hypersense, yun2025missionhd]. A conventional HDC classifier stores one prototype per class and predicts by comparing a query $\phi(\mathbf{x})$ to all $C$ prototypes [hernandez2021onlinehd]. The resulting memory cost $\mathcal{O}(CD)$ dominates in multi-class, resource-constrained regimes, so shrinking this footprint without eroding accuracy or robustness is central.
Prior compression largely targets the feature axis by reducing $D$ [imani2019sparsehd, morris2019comphd, ge2020classification, verges2025classification]. As a representative state-of-the-art example, SparseHD [imani2019sparsehd] sparsifies trained hypervectors to accelerate computation while keeping one prototype per class. However, operating at lower effective dimensionality weakens robustness to hardware noise—an effect consistent with observations from memory-centric and analog/near-memory accelerators [burr2023design, frank2023impact, lammie2025deep, krestinskaya2024neural] as illustrated in Figure 1.(a).
@@
We propose LogHD, a class-axis compression scheme that replaces the $C$ per-class prototypes with $n$ bundle hypervectors $\{\mathbf{M}_{j}\}_{j=1}^{n}$, where $n\!\geq\!\lceil\log_{k}C\rceil$ for an alphabet size $k\!\geq\!2$. Each class receives a length-$n$ $k$-ary code that specifies how its prototype contributes to each bundle. At inference, a query is compared to the $n$ bundles to form an activation vector that is decoded via learned per-class profiles. This reduces memory from $\mathcal{O}(CD)$ to $\mathcal{O}(nD)=\mathcal{O}(D\log_{k}C)$ without shrinking $D$. The approach is orthogonal to feature-axis methods (including SparseHD-style sparsification) and aligns with emerging HDC/VSA macros and cross-layer flows that benefit from storing fewer vectors rather than shorter ones [nayan2025hydra, du2025cross] as depicted in Figure 1.(b).
Preserving $D$ has direct implications for robustness. In noisy memories and in-/near-memory fabrics, random bit flips or small analog perturbations distort stored hypervectors and degrade similarity computations [yun2023hyperdimensional, barkam2023reliable]. Concentration-of-measure effects stabilize similarities at large $D$, whereas reducing $D$ amplifies score variance. By decreasing the number of stored hypervectors instead of their length, LogHD achieves compactness while retaining high-$D$ tolerance to non-idealities, consistent with reliability analyses of analog and near-memory accelerators [burr2023design, frank2023impact, tsai2024architecture]. Orthogonal techniques—such as flexible numeric types and dynamic dimensional masking—can be layered with LogHD when tighter budgets are required [li2025fate, liu2025hyperdyn]. Thus, we also study a hybrid design that combines LogHD with SparseHD-style sparsification, offering additional memory savings with robustness between pure LogHD and aggressive feature-axis compression as indicated in Figure 1.(c).
@@ II-A Foundations of HDC
Hyperdimensional computing (also known as vector–symbolic architectures) represents symbols and structures as dense, high-dimensional hypervectors and manipulates them via binding, bundling, and permutation [kanerva2009hyperdimensional]. In classification, training samples for a class are superposed into a prototype hypervector; inference compares a query to all class prototypes using cosine or Hamming similarity [hernandez2021onlinehd]. The paradigm maps well to in-/near-memory accelerators because its operations are linear and massively parallel, and it exhibits empirical resilience to device non-idealities [yun2023hyperdimensional, barkam2023reliable, barkam2023hdgim]. Comparative studies further catalog encoding/training choices and robustness trends, situating HDC among compact, hardware-friendly learners [verges2025classification, zou2021scalable, wang2023disthd].
@@ II-B Model-size reduction in HDC
Most prior work reduces parameters along the feature axis while retaining the one prototype per class layout. QuantHD focuses on precision reduction (binary/ternary encodings) so that Hamming distance replaces cosine similarity, improving memory footprint and compare cost without changing the number of stored prototypes [imani2019quanthd, verges2025classification]. CompHD packs information within each prototype via structured splitting and binding with positional hypervectors followed by superposition; the query path mirrors these transforms at inference [morris2019comphd]. This yields a complex feature-axis reduction that preserves per-class prototypes and introduces additional binding/unbinding overhead and cross-talk considerations [verges2025classification]. SparseHD sparsifies trained prototypes and co-optimizes the algorithm with reconfigurable hardware, achieving state-of-the-art efficiency–accuracy trade-offs among few-parameter compression methods while still storing one (sparsified) prototype per class [imani2019sparsehd, verges2025classification]. In our study, SparseHD is therefore used as the representative feature-axis compressor—both as a strong baseline for performance/robustness comparisons and as the feature-axis component in our hybrid LogHD+SparseHD design—whereas CompHD is not used for hybridization due to its more complex binding/unbinding pipeline and sensitivity to cross-talk. Beyond these, accuracy-oriented updates such as OnlineHD and distributed/regenerative variants improve learning dynamics but leave the parameter layout unchanged, i.e., they do not reduce the number of class prototypes [hernandez2021onlinehd, zou2021scalable, wang2023disthd, verges2025classification]. In contrast, our approach compresses along the class axis: rather than shrinking each prototype, it replaces the $C$ prototypes by $n\!\approx\!\lceil\log_{k}C\rceil$ bundles and decodes in the induced activation space. To the best of our knowledge and as reflected in recent surveys [verges2025classification], log-scale reduction in the number of stored class hypervectors has not been explored in HDC.
@@ II-C Robustness to hardware noise
HDC’s robustness is rooted in concentration at high dimensionality: for a given bit-flip or analog error rate, larger $D$ averages perturbations and stabilizes similarity scores, whereas reducing $D$ increases score variance [yun2023hyperdimensional, barkam2023reliable]. Feature-axis compression (CompHD, SparseHD) therefore trades some of this averaging for footprint. By compressing along the class axis and preserving $D$, our method maintains the dimensionality-driven tolerance to non-idealities while still reducing overall model size, a behavior consistent with reliability observations reported for memory-centric ML on in-/near-memory substrates [verges2025classification].
III Methodology
@@ III-A Preliminaries
$\delta(\mathbf{u},\mathbf{v})=\Big\langle\tfrac{\mathbf{u}}{\|\mathbf{u}\|_{2}},\,\tfrac{\mathbf{v}}{\|\mathbf{v}\|_{2}}\Big\rangle,$
@@ III-D Generating Activation Profiles
$A(\mathbf{x})\;=\;\big(\delta(\mathbf{M}_{1},\phi(\mathbf{x})),\,\dots,\,\delta(\mathbf{M}_{n},\phi(\mathbf{x}))\big)\in\mathbb{R}^{n}.$
@@
$\mathbf{P}_{y}\;=\;\mathbb{E}_{\mathbf{x}\mid y}\big[A(\mathbf{x})\big]\;\approx\;\frac{1}{N_{y}}\sum_{i=1}^{N_{y}}A\!\big(\mathbf{x}^{(y)}_{i}\big)$,
@@ IV Experiments
ISOLET [isolet_54] 617 26 6,238 1,559 Voice recognition
UCIHAR [human_activity_recognition_using_smartphones_240] 261 12 6,213 1,554 Activity recognition (mobile)
PAMAP2 [pamap2_physical_activity_monitoring_231] 75 5 611,142 101,582 Activity recognition (IMU)
PAGE [page_blocks_classification_78] 10 5 4,925 548 Page layout blocks classification
@@ V Conclusions
We presented LogHD, a class-axis compression scheme for hyperdimensional computing that reduces memory and compute from $\mathcal{O}(CD)$ to $\mathcal{O}(D\log_{k}C)$ while preserving dimensionality. By storing fewer hypervectors rather than shorter ones, LogHD achieves robustness to device faults and quantization, and in hybrid form offers additional flexibility under tight budgets. Experiments across multiple datasets show that LogHD sustains accuracy under up to $3\times$ higher bit-flip rates than feature-axis compression, and ASIC implementations deliver up to $498\times$ energy efficiency and $63\times$ speedup over CPU baselines. These results demonstrate that class-axis compaction is a scalable and hardware-friendly strategy for robust memory-centric machine learning.
@@ Acknowledgements
This work was supported in part by the DARPA Young Faculty Award, the National Science Foundation (NSF) under Grants #2431561, #2127780, #2319198, #2321840, #2312517, and #2235472, the Semiconductor Research Corporation (SRC), the Office of Naval Research through the Young Investigator Program Award and Grants #N00014-21-1-2225 and #N00014-22-1-2067, Army Research Office Grant #W911NF2410360, the U.S. Army Combat Capabilities Development Command Army Research Laboratory under Support Agreement No. USMA 21050, and DARPA under Support Agreement No. USMA 23004. Additionally, support was provided by the Air Force Office of Scientific Research under Award #FA9550-22-1-0253, along with generous gifts from Xilinx and Cisco.

### v2

@@ I Introduction
Machine-learning architectures operating under tight memory, energy, and reliability budgets increasingly depend on cross-layer co-design and in-/near-memory substrates. Recent surveys and system studies underline both the promise and the sensitivity of memory-centric ML to device non-idealities and limited precision, motivating methods that reduce footprint while preserving robustness [14, 13, 22, 3, 6]. Within this context, hyperdimensional computing (HDC)—a family of vector–symbolic architectures (VSA)—is a strong fit: its operations are inherently parallel and memory-centric, and recent cross-layer workflows and macro-level prototypes indicate practical routes to efficient HDC/VSA hardware in embedded and in-/near-memory settings [19, 5].
HDC encodes data as $D$-dimensional hypervectors and learns via light-weight algebraic operations that are parallel, memory-centric, and empirically tolerant to device noise [12, 29, 10, 30, 24, 28, 25, 26, 27]. A conventional HDC classifier stores one prototype per class and predicts by comparing a query $\phi(\mathbf{x})$ to all $C$ prototypes [8]. The resulting memory cost $\mathcal{O}(CD)$ dominates in multi-class, resource-constrained regimes, so shrinking this footprint without eroding accuracy or robustness is central.
Prior compression largely targets the feature axis by reducing $D$ [11, 18, 7, 23]. As a representative state-of-the-art example, SparseHD [11] sparsifies trained hypervectors to accelerate computation while keeping one prototype per class. However, operating at lower effective dimensionality weakens robustness to hardware noise—an effect consistent with observations from memory-centric and analog/near-memory accelerators [3, 6, 14, 13] as illustrated in Figure 1.(a).
@@
We propose LogHD, a class-axis compression scheme that replaces the $C$ per-class prototypes with $n$ bundle hypervectors $\{\mathbf{M}_{j}\}_{j=1}^{n}$, where $n\!\geq\!\lceil\log_{k}C\rceil$ for an alphabet size $k\!\geq\!2$. Each class receives a length-$n$ $k$-ary code that specifies how its prototype contributes to each bundle. At inference, a query is compared to the $n$ bundles to form an activation vector that is decoded via learned per-class profiles. This reduces memory from $\mathcal{O}(CD)$ to $\mathcal{O}(nD)=\mathcal{O}(D\log_{k}C)$ without shrinking $D$. The approach is orthogonal to feature-axis methods (including SparseHD-style sparsification) and aligns with emerging HDC/VSA macros and cross-layer flows that benefit from storing fewer vectors rather than shorter ones [19, 5] as depicted in Figure 1.(b).
Preserving $D$ has direct implications for robustness. In noisy memories and in-/near-memory fabrics, random bit flips or small analog perturbations distort stored hypervectors and degrade similarity computations [25, 1]. Concentration-of-measure effects stabilize similarities at large $D$, whereas reducing $D$ amplifies score variance. By decreasing the number of stored hypervectors instead of their length, LogHD achieves compactness while retaining high-$D$ tolerance to non-idealities, consistent with reliability analyses of analog and near-memory accelerators [3, 6, 22]. Orthogonal techniques—such as flexible numeric types and dynamic dimensional masking—can be layered with LogHD when tighter budgets are required [15, 16]. Thus, we also study a hybrid design that combines LogHD with SparseHD-style sparsification, offering additional memory savings with robustness between pure LogHD and aggressive feature-axis compression as indicated in Figure 1.(c).
@@ II-A Foundations of HDC
Hyperdimensional computing (also known as vector–symbolic architectures) represents symbols and structures as dense, high-dimensional hypervectors and manipulates them via binding, bundling, and permutation [12]. In classification, training samples for a class are superposed into a prototype hypervector; inference compares a query to all class prototypes using cosine or Hamming similarity [8]. The paradigm maps well to in-/near-memory accelerators because its operations are linear and massively parallel, and it exhibits empirical resilience to device non-idealities [25, 1, 2]. Comparative studies further catalog encoding/training choices and robustness trends, situating HDC among compact, hardware-friendly learners [23, 30, 24].
@@ II-B Model-size reduction in HDC
Most prior work reduces parameters along the feature axis while retaining the one prototype per class layout. QuantHD focuses on precision reduction (binary/ternary encodings) so that Hamming distance replaces cosine similarity, improving memory footprint and compare cost without changing the number of stored prototypes [9, 23]. CompHD packs information within each prototype via structured splitting and binding with positional hypervectors followed by superposition; the query path mirrors these transforms at inference [18]. This yields a complex feature-axis reduction that preserves per-class prototypes and introduces additional binding/unbinding overhead and cross-talk considerations [23]. SparseHD sparsifies trained prototypes and co-optimizes the algorithm with reconfigurable hardware, achieving state-of-the-art efficiency–accuracy trade-offs among few-parameter compression methods while still storing one (sparsified) prototype per class [11, 23]. In our study, SparseHD is therefore used as the representative feature-axis compressor—both as a strong baseline for performance/robustness comparisons and as the feature-axis component in our hybrid LogHD+SparseHD design—whereas CompHD is not used for hybridization due to its more complex binding/unbinding pipeline and sensitivity to cross-talk. Beyond these, accuracy-oriented updates such as OnlineHD and distributed/regenerative variants improve learning dynamics but leave the parameter layout unchanged, i.e., they do not reduce the number of class prototypes [8, 30, 24, 23]. In contrast, our approach compresses along the class axis: rather than shrinking each prototype, it replaces the $C$ prototypes by $n\!\approx\!\lceil\log_{k}C\rceil$ bundles and decodes in the induced activation space. To the best of our knowledge and as reflected in recent surveys [23], log-scale reduction in the number of stored class hypervectors has not been explored in HDC.
@@ II-C Robustness to hardware noise
HDC’s robustness is rooted in concentration at high dimensionality: for a given bit-flip or analog error rate, larger $D$ averages perturbations and stabilizes similarity scores, whereas reducing $D$ increases score variance [25, 1]. Feature-axis compression (CompHD, SparseHD) therefore trades some of this averaging for footprint. By compressing along the class axis and preserving $D$, our method maintains the dimensionality-driven tolerance to non-idealities while still reducing overall model size, a behavior consistent with reliability observations reported for memory-centric ML on in-/near-memory substrates [23].
@@ III-A Preliminaries
$\delta(\mathbf{u},\mathbf{v})=\Big\langle\tfrac{\mathbf{u}}{\|\mathbf{u}\|_{2}},\,\tfrac{\mathbf{v}}{\|\mathbf{v}\|_{2}}\Big\rangle,$ (1)
@@ III-D Generating Activation Profiles
$A(\mathbf{x})\;=\;\big(\delta(\mathbf{M}_{1},\phi(\mathbf{x})),\,\dots,\,\delta(\mathbf{M}_{n},\phi(\mathbf{x}))\big)\in\mathbb{R}^{n}.$ (5)
@@
$\mathbf{P}_{y}\;=\;\mathbb{E}_{\mathbf{x}\mid y}\big[A(\mathbf{x})\big]\;\approx\;\frac{1}{N_{y}}\sum_{i=1}^{N_{y}}A\!\big(\mathbf{x}^{(y)}_{i}\big)$, (6)
@@ III-H Practical Considerations
ISOLET [4] 617 26 6,238 1,559 Voice recognition
UCIHAR [21] 261 12 6,213 1,554 Activity recognition (mobile)
PAMAP2 [20] 75 5 611,142 101,582 Activity recognition (IMU)
PAGE [17] 10 5 4,925 548 Page layout blocks classification
IV Experiments
@@ V Conclusions
We presented LogHD, a class-axis compression scheme for hyperdimensional computing that reduces memory and compute from $\mathcal{O}(CD)$ to $\mathcal{O}(D\log_{k}C)$ while preserving dimensionality. By storing fewer hypervectors rather than shorter ones, LogHD achieves robustness to device faults and quantization, and in hybrid form offers additional flexibility under tight budgets. Experiments across multiple datasets show that LogHD sustains accuracy under up to $3\times$ higher bit-flip rates than feature-axis compression, and ASIC implementations deliver up to $498\times$ energy efficiency and $63\times$ speedup over CPU baselines. Together, these results position class-axis compaction as a scalable, hardware-friendly approach for noise-robust HDC model compression, well suited for TinyML and edge deployments.
Limitations and Future Work. This work does not include a formal theoretical analysis of how class-axis reduction interacts with hypervector capacity and separability as dimensionality and class complexity grow. Exploring these relationships, along with richer fault models, is left to future work.
@@ Acknowledgements
This work was supported in part by the DARPA Young Faculty Award, the National Science Foundation (NSF) under Grants #2431561, #2127780, #2319198, #2321840, #2312517, and #2235472, the Semiconductor Research Corporation (SRC), the Office of Naval Research through the Young Investigator Program Award and Grants #N00014-21-1-2225 and #N00014-22-1-2067, Army Research Office Grant #W911NF2410360, and DARPA under Support Agreement No. USMA 23004. Additionally, support was provided by the Air Force Office of Scientific Research under Award #FA9550-22-1-0253, along with generous gifts from Xilinx and Cisco.